Homework 1 11-721: Grammars and Lexicons, Fall 2008 Due Monday, September 15, 2008

1 Predeterminers, Determiners, and Adjectives

In this problem you will examine the distributional and morphological properties of three parts of speech in English: predeterminers, determiners, and adjectives. This is a realistic exercise: you are building a lexicon for your machine translation system or question answering system, and you need to decide on the parts of speech for some words. You look in a book and you find some tests for distinguishing determiners and adjectives. You have to apply the tests, imperfect as they are, and make your decisions.

The exercise has several purposes:

1. You will learn to apply tests.

2. You will learn to interpret the results of tests.

3. You will encounter problems in categorizing words into parts of speech.

1.1 The Tests: distributional and morphological properties of predeterminers, determiners, and adjectives in English

1.1.1 Determiner Test 1:

Test: Determiners do not co-occur with the and a.

Example: Is the word this a determiner?

Test Sentences:

(1) a.* The this book is about linguistics.

b.* This the book is about linguistics.

c.* A this book is about linguistics.

d.* This a book is about linguistics.

Conclusion: The word this passes Test 1 for determiners.

1.1.2 Determiner Test 2: Test: If a determiner can occur with singular count nouns, it makes them acceptable.

You probably need to know some background in order to understand this test. English has count nouns and mass nouns, as do many other languages. Count nouns can occur with cardinal numbers or with a. Mass nouns can occur with some. Count nouns:

(2) a. Two books are on the table. b. A book is on the table. c. Twenty students are in the class. d. A student is in the lounge.

Mass nouns:

(3) a. Some sand is on the table. b.* A sand is on the table. c.* Two sands are on the table. d. Some information is available on the Web. e.* An information is available on the Web. f.* Two informations are available on the Web.

The next fact is critical for Test 2: Singular count nouns cannot occur without a determiner in English:

(4) a.* Book is on the table.

b.* I saw book. c.* Student studied. d.* I talked to student.

Singular count nouns are fine with a determiner:

(5) a. The/a book is on the table. b. I saw the/a book. c. The/a student studied. d. I talked to the/a student.

Example: Is the word this a determiner?

Test Sentences:

(6) a. This book is on the table.

b. I saw this book.

c. This student studied. d. I talked to this student.

Conclusion: This passes Test 2 for determiners.

Example: Is the word these a determiner?

Conclusion: Test 2 for determiners is not applicable to these because these only occurs with plural nouns. 1.1.3 Determiner Test 3: Test: Determiners that are about quantities can participate in a partitive construction.

In Examples (7)a and b, the partitive construction is each of the students. It is called partitive because it is talking about part of the set of students.

(7) a. Each of the students has studied.

b. I talked to each of the students. This test is not applicable to a, the, this, that, these, and those because they are not about quantities. They express definiteness and proximity. 1.1.4 Predeterminer Test: Test: There is just one test for predeterminers. They can occur before a or the.

(8) a. What a party it was!

b. I never saw so big an elephant. c. All the students left.

1.1.5 Adjective Test 1: Test: Most adjectives can occur between the or a and a noun.

(9) a. The blue book is on the table. b. A blue book is on the table. c. I saw a blue book. d. I saw the blue book.

1.1.6 Adjective Test 2: Test: Most adjectives can occur in predicative position, after verbs like be, become, seem, or look.

(10) a. The test seemed difficult. b. The book was blue. c. The cut became infected. d. The cake looked yummy.

Some adjectives can only occur in predicative position:

(11) a.* I saw an asleep child. b. The child was asleep. c.* The ahead runner won the race. d. The runner was ahead.

1.1.7 Adjective Test 3: Test: Adjectives that describe gradable properties have comparative and superlative forms:

(12) a. This copy is darker.

b. The darker copy is better. c. The darkest copy is not good. d. This student is more studious. e. The most studious student won a prize.

1.1.8 Adjective Test 4:

Test: Adjectives that describe gradable properties can be modified with very.

(13) a. That book is very blue.

b. The very light print is not readable.

1.1.9 Adjective Test 5:

Test: Adjectives that describe gradable properties can be modified by so in predicative positions, but not pre-nominal positions:

(14) a. That book is so blue.

b. The sky seemed so dark.

(15) a.* I saw the so blue book.

b.* I saw so blue water.

c.* The so light print is not readable.

Don’t get confused by this: an adjective with so cannot occur between a determiner and a noun, but it can occur as a predeterminer, before the determiner:

(16) I never saw so big an elephant.

1.1.10 Determiner Test 4:

Test: A determiner that describes a gradable property may be modified with so. In order to pass this test, the word must be in pre-nominal position.

(17) He made so many mistakes.

Determiner Test 4 and Adjective Test 5 are easily confusable, and in fact, some pieces of data might be relevant to both tests. In order to pass Adjective Test 5, the word must be able to appear with so in predicative position, but not in prenominal position. In order to pass Determiner Test 4, the word must be gradable and occur with so in prenominal position.

1.1.11 Adjective Test 6:

Test: Another partitive construction: Gradable adjectives may occur in this construction.

(18) a. The largest of the books couldn’t fit in the box.

b. I have addressed the most important of your criticisms.

Determiner Test 3 and Adjective Test 6 are easily confused. Determiner Test 3 is about partitive constructions like many of the students. Adjective Test 6 is about a construction that involves a superlative plus a partitive (the largest of the boxes).

1.2 Summary of tests

• Predeterminer Test: Occurs before the, a, or an.

Template: PREDETERMINER the/a/an NOUN

• Determiner Test 1: Does not occur with the, a, or an.

Template: * DETERMINER the/a/an NOUN
Template: * the/a/an DETERMINER NOUN

• Determiner Test 2: Makes a singular count noun acceptable.

Template: DETERMINER COUNT-NOUN

• Determiner Test 3: Partitive construction.

Template: DETERMINER of the NOUN

• Determiner Test 4: Gradable determiners occur with so pre-nominally.

Template: so GRADABLE-DETERMINER NOUN

• Adjective Test 1: Occurs between the, a, or an and a noun.

Template: the/a/an ADJECTIVE NOUN

• Adjective Test 2: Occurs in predicative position.

Template: be/seem/become/look ADJECTIVE

• Adjective Test 3: Gradable adjectives have comparative and superlative forms.

Template: more ADJECTIVE
Template: most ADJECTIVE
Template: ADJECTIVE-er
Template: more ADJECTIVE-est

• Adjective Test 4: Gradable adjectives can be modified by very.

Template: very ADJECTIVE

• Adjective Test 5: Gradable adjectives do not occur prenominally with so, but do occur in predicative position with so.

Template: * DETERMINER (optional) so ADJECTIVE NOUN Template:
be/seem/become/look so ADJECTIVE

• Adjective Test 6: Gradable adjectives occur in the partitive superlative construction.

Template: the ADJECTIVE-est of the NOUN Template: the most ADJECTIVE
of the NOUN

1.3 Interpreting the results of tests

Write two things in each cell of the following table: (1) P if the word passed the test or F if the word failed and (2) the number of the example that supports your conclusion. Leave the cell blank if there was no relevant data. Use my grammaticality judgments.

There are a few possibilities:

• There was no relevant data: leave the cell blank.

• There is one piece of relevant data: put the example number and the test result (P or F) in the cell.

• There is more than one piece of relevant data, and both have the same result. We will count it right if you list one example number in the cell.

• There is more than one piece of relevant data, and the results are different. This might or might not arise depending on how you interpret the tests. If you think you have found a case like this, write both results and both example numbers in the cell.

enough sufficient much many

PreDet Det1 Det3 Det4 Adj1 Adj3 Adj4 Adj5 Adj6

(19) a.* Enough the books were given to the class.

b.* The schoolboard provided the enough books.

c. The student has read enough of the books.

d.* The bank gave them so enough money.

e.* The funds were so enough.

f.* The funds were very enough.

g.* The most enough of the funds.

(20) a. The accounts were filled with the sufficient funds. b. The banker gave them a sufficient amount. c.* Sufficient of the books were given to the class. d. These funds were very/so/more sufficient. e. The most sufficient of the funds.

(21) a.*Much the/a water was spilled. b.*The much water was spilled. c. Much of the water was spilled. d. So much water was spilled. e. The water was so much.

(22) a. Many a student has taken this class. b. I remember the many students who have taken this class. c. Many of the students have taken this class. d. I saw very many students. e. So many students have taken this class. f. The problems were so many.

1.4 Applying the tests

• Apply each test to the word numerous. You should write one or two

sentences for each test. For each test, you should do three things:

– Apply the test: Insert the word numerous into each template mechanically and write down a sentence or phrase.

– Judge grammaticality: Judge whether the sentence or phrase is grammatical. If you are not a native speaker of English, ask a native speaker for a grammaticality judgment. Then interpret the results.

– Interpret the results: Did numerous pass or fail the test?

• Find one test for determiners that each passes but every does not. Write down two sentences or phrases, one showing that each passes the test and one showing that every does not pass it.